\begin{thebibliography}{77}

% \bibitem{Bab} 
% K. I. Babenko, 
% {\it About the approximation of periodic functions of many variable trigonometric polynomials},
% Doklady Akademii Nauk SSSR, 32, 247--250, 1960.

\bibitem{Bauhardt}
W.~Bauhardt, \textit{Hilbert-Zahlen von Operatoren in Banachr\"aumen}, Math. Nachr. 79, 
181--187, 
% pp.~181, 
1977.

% \bibitem{BSS} J.D. Batson, D.A. Spielman, N. Srivastava, 
% {\it Twice-Ramanujan sparsifiers}, in: STOC’09—Proceedings of the 2009 ACM International Symposium on Theory of Computing, pp. 255–-262, ACM, New York, 2009.


% \bibitem{Bernardi_Maday1992}
% C. Bernardi and Y. Maday, 
% {\it  Polynomial interpolation results in {S}obolev spaces},
% {J.~Comput. Appl. Math.},
%  {43(1)}, 53--80,
%  {1992},
%     {https://doi.org/10.1016/0377-0427(92)90259-Z}.

% \bibitem{Binev_Bonito_2023}
% P. Binev, A. Bonito, A. Cohen, W. Dahmen, R. DeVore, and G. Petrova,
% {\it Solving PDEs with Incomplete Information}, arXiv:2301.05540. 

% \bibitem{Bos18}
% L.~Bos, 
% \textit{Fekete points as norming sets}, 
% Dolomites Research Notes on Approximation 11(4), 26--34, 2018. 
% %doi: 10.14658/pupj-drna-2018-4-3

%\bibitem{Bugrov1964} 
%S. Bugrov,
%{\it Approximation of a class of functions with a dominant mixed derivative}. 
%Mat. Sb., 64(106), 410--418, 1964.


% \bibitem{Byrenheid_Dung_Sickel_Ullrich}
% G.\,Byrenheid, D.\,D\~ung, W.\,Sickel, and T.\,Ullrich,
% {\it Sampling on energy-norm based sparse grids for the optimal recovery of {S}obolev type functions in $H^\omega$},
% J. Approx. Theory,
%  207, 207--231,
% 2016,
% https://doi.org/10.1016/j.jat.2016.02.012.

% \bibitem{Byrenheid_Ullrich2017} G. Byrenheid and T. Ullrich,
% {\it Optimal sampling recovery of mixed order Sobolev embeddings
% via discrete Littlewood-Paley type characterizations}, Analisis Matematica, 2017.

% \bibitem{CL08}
% J.-P. Calvi, N. Levenberg, 
% Uniform approximation by discrete least squares polynomials, 
% J. Approx. Theory 152(1), 82--100, 2008.

% \bibitem{Carl81}
% B. Carl, \textit{Entropy numbers, s-numbers, and eigenvalue problems}, J. Functional Analysis 41(3), 290--306, 1981.

% \bibitem{CP77}
% B. Carl, A. Pietsch, \textit{Entropy numbers of operators in Banach spaces}, 
% In: Novák, J. (eds) General Topology and Its Relations to Modern Analysis and Algebra IV. Lecture Notes in Mathematics, vol. 609, Springer, Berlin, Heidelberg, 1977.

% \bibitem{CD23}
% A. Chkifa and M. Dolbeault, 
% {\it Randomized least-squares with minimal oversampling and
% interpolation in general spaces}, 
% arXiv:2306.07435.

%\bibitem{CKS} F. Cobos, T. Kühn, and  W. Sickel, 
%{\it Optimal approximation of multivariate periodic
%Sobolev functions in the sup-norm},
%Journal of Functional Analysis, 270(11), 4196--4212, 2016.

%\bibitem{CD} A. Cohen, M. Dolbeault, {\it Optimal pointwise sampling for $L_2$ approximation}, Journal of Complexity, 68, 101602, 2022.

% \bibitem{CM} A. Cohen, G. Migliorati, 
% \emph{Optimal weighted least squares methods,} SMAI J. Comput. Math., 3, 181-203, 2017.

% \bibitem{Creutzig_Wojtaszczyk2004} 
% J. Creutzig, P. Wojtaszczyk,
% {\it Linear vs. nonlinear algorithms for linear problems},
% J. Complexity, 20, 807--820, 2004. 
% %doi:10.1016/j.jco.2004.05.003

% \bibitem{DPTT19}
% F. Dai, A. Prymak, V.N. Temlyakov, S.Yu. Tikhonov, 
% {\it Integral norm discretization and related problems}, Russ. Math.
% Surv. 74 (4) (2019) 579--630
% %, arXiv:1807.01353v1
% . Translation from Usp. Mat. Nauk 74:4 (448), 3--58, 2019.

% \bibitem{Diestel-book}
% J. Diestel, 
% {\it Geometry of Banach Spaces -- Selected Topics}, 
% Lecture Notes in Mathematics, Vol. 485. Springer-Verlag, Berlin-New York, 
% 1975, xi+282 pp.

% \bibitem{DKU} M.\,Dolbeault, D.\,Krieg, and M.\,Ullrich,  
% {\it A sharp upper bound for sampling numbers in $L_2$}, 
% {Appl. Comput. Harmon. Anal.} 63, 113--134, 2023. 
% %		\href{https://doi.org/10.1016/j.acha.2022.12.001}{DOI:10.1016/j.acha.2022.12.001}. 
% %arXiv:2204.12621, 2022.


% \bibitem{DD91} D.\,D{\~u}ng, {\it On optimal recovery of multivariate periodic functions}, In: Har-
% monic Analysis (Conference Proceedings, Ed. S. Igary), Springer-Verlag
% 1991, Tokyo-Berlin.


%\bibitem{DTU18}
%{D.}~D\~ung, V.\,N.~Temlyakov, and T.~Ullrich,
%\newblock {H}yperbolic {C}ross {A}pproximation.
%\newblock {\em Advanced Courses in Mathematics - CRM Barcelona}.
%  Springer International Publishing, 2018.

\bibitem{FMS21}
 D. L. Fernandez, M. Masty\l o, and E. B. Silva. \textit{Pietsch’s variants of s -numbers for multilinear operators}, 
Rev. R. Acad. Cienc. Exactas Fís. Nat. Ser. A Mat. RACSAM 115(4):Paper No. 189, 24, 2021.

% \bibitem{FHJU22} F. Filbir, R. Hielscher, T. Jahn, and T. Ullrich, {\it Marcinkiewicz--Zygmund inequalities for scattered and random data on the $q$-sphere}, arXiv:2303.00045. 


% \bibitem{Freud} G. Freud,
% {\it Orthogonale Polynome}. Lehrbücher und Monographien aus dem Gebiete
% der Exakten Wissenschaften, Mathematische Reihe, Band 33. Birkh\"auser Verlag, Basel-Stuttgart, 1969.

%\bibitem{Galeev1978} E. M. Galeev,
%{\it Approximation of classes of functions with several bounded
%derivatives by Fourier sums}.
%\newblock {\em Mat. Zametki}, 23, 197--212, 1978; English
%transl. in Mathematical Notes 23, 1978.

% \bibitem{GePe14}
% D.~Geller, I.Z.~Pesensen,
% \emph{Komogorov and linear width of balls in Sobolev spaces on compact manifolds}, 
% Math. Scand. (2014), Vol. 115, No. 1, pp. 96--122.

% \bibitem{GW23}
% J.~Geng and H.~Wang, 
% {\it On the power of standard information for tractability for $L_\infty$ approximation of periodic functions in the worst case setting}, 
% arXiv:2304.14748.

% \bibitem{Gin}
% E.~Gin\'e, {\it The addition formula for the eigenfunctions of the Laplacian},
% Advances in Mathematics, 18(1):102--107, 1975.


% \bibitem{Griebel_Knapek2000}
% M.\,Griebel and S.\,Knapek, 
% {\it Optimized tensor-product approximation spaces}, 
% Constr. Approx., 16(4), 525--540, 2000, https://doi.org/10.1007/s003650010010.

% \bibitem{Griebel_Knapek2009} 
% M.\,Griebel and S.\,Knapek, 
% {\it Optimized general sparse grid approximation spaces for operator equations},
% Math. Comp.,
% 78(268), 2223--2257, 2009, https://doi.org/10.1090/S0025-5718-09-02248-0.

% \bibitem{Gr19} K. Gröchenig, {\it Sampling, Marcinkiewicz-Zygmund inequalities, approximation, and quadrature rules}, J. Approx. Theory, 257:105455, 2020.

% \bibitem{Heinrich} S. Heinrich, {\it Random approximation in numerical analysis}, in K. D. Bierstedt and et al.,
% editors, Functional Analysis, pages 123–171. Dekker, New York, 1994.
	
 % \bibitem{HNV} A.\,Hinrichs, E.\,Novak, and J.\,Vybiral, {\it  Linear information versus function evaluations for $L_2$-approximation}, J. Approx. Theory, 153(1),  97-107, 2008.

%\bibitem{HKNPU} A. Hinrichs, D. Krieg, E. Novak, J. Prochno and M. Ullrich, 
%{\it Random sections of ellipsoids and the power of random information},
%Transactions of the American Mathematical Society 374(12), 8691-8713, 2021.

% \bibitem{JL-handbook}
% W. B. Johnson, J. Lindenstrauss (eds.), \textit{Handbook of the Geometry of Banach Spaces}, 1. ed., Elsevier, 2001.

\bibitem{HL84} 
S. Heinrich, R. Linde, 
\textit{On the asymptotic behaviour of Hilbert numbers}, 
Math. Nachr. 119, 117--120, 1984.

%\bibitem{HKNPU2}
%A.~Hinrichs, D.~Krieg, E.~Novak, J.~Prochno, and M.~Ullrich.
%\newblock On the power of random information.
%\newblock In F.~J. Hickernell and P.~Kritzer, editors, {\em Multivariate Algorithms and Information-Based Complexity}, 43-64. De Gruyter, Berlin/Boston, 2020.

%\bibitem{HKNVa}
%A. Hinrichs, D. Krieg, E. Novak and J. Vybiral, 
%{\it Lower bounds for the error of quadrature formulas for Hilbert spaces},
%Journal of Complexity, 65, 101544, 2021, https://doi.org/10.1016/j.jco.2020.101544.


% \bibitem{HKNV} A.\,Hinrichs, D.\,Krieg, E.\,Novak, and J.\,Vybiral, 
% {\it Lower bounds for integration and recovery in $ L_2$}, J. Complexity, 101662, 2022. % https://doi.org/10.1016/j.jco.2022.101662.

% \bibitem{HM} S. Hubbert and T.\,M.~Morton,
% {\it $L_p$-error estimates for radial basis function interpolation on the sphere},
% J. Approx. Theory, 129(1):58--77, 2004.

\bibitem{Ism74}
R. S. Ismagilov, 
\textit{Diameters of sets in normed linear spaces and approximation of functions by trigonometric polynomials}, Russian Math. Surveys 29(3), 169--186, 1974. 

% \bibitem{JUV} T. Jahn, T. Ullrich, F. Voigtlaender,
% {\it Sampling numbers of smoothness classes via $\ell^1$-minimization},
% J. Complexity 79, 101786, 2023.

%\bibitem{KUV} L. K\"ammerer, T. Ullrich, and T. Volkmer, 
%{\it Worst case recovery guarantees for least squares approximation using random samples}, 
%Constr. Approx., 2021, https://doi.org/10.1007/s00365-021-09555-0.
%arXiv: 1911.10111, 2019, to appear in Constructive Approximation.

% \bibitem{KS} M.\,I.~Kadets and M.\,G.~Snobar,
% %{\it Certain functionals on the Minkowski compactum}, Mathematical Notes, 10:694--696, 1971.
% {\it Some functionals over a compact Minkovskii space}, Math. Notes, 10(4), 694--696, 1971.
% https://link.springer.com/article/10.1007/BF01106467

% \bibitem{Kashin77} 
% B.\,S.\,Kashin, 
% {\it Diameters of some finite-dimensional sets and
% classes of smooth functions}, 
% Izv. Akad. Nauk SSSR Ser. Mat., 41(2): 334--351, 1977; English transl. in Math. USSR-Izv.,
% 11(2), 317--333, 1977,
% https://doi.org/10.1070/IM1977v011n02ABEH001719.

% \bibitem{KKT21}
% B.~Kashin, S.~Konyagin, and V.~Temlyakov, 
% {\it Sampling discretization of the uniform norm},
% Constr. Approx. 57(2), 663--694, 2023.

% \bibitem{KKLT} B. Kashin, E. Kosov, I. Limonova, and V. Temlyakov,
% {\it Sampling discretization and related problems},
% J. Complexity, 101653, 2022.
% %https://doi.org/10.1016/j.jco.2022.101653.

% \bibitem{Kol_Lomako_Tikhonov23}
% Y.\,Kolomoitsev, T.\,Lomako, and S.\,Tikhonov, 
% {\it Sparse grid approximation in weighted {W}iener spaces},
% J. Fourier Anal. Appl.,  29 (19), 2023, https://doi.org/10.1007/s00041-023-09994-2.

\bibitem{Koe86}
H. K\"onig, {\it Eigenvalue distribution of compact operators}, 
Operator Theory: Advances and Applications 16, Birkhäuser, Basel, 1986.

% \bibitem{Koe01}
% H. K\"onig, {\it Eigenvalues of operators and applications} 
% In Handbook of the geometry of Banach spaces, Vol. I, North-Holland, Amsterdam, 2001, 941--974.

% \bibitem{Kam19} L. K\"ammerer,
% {\it Multiple lattice rules for multivariate {$L_{\infty}$} approximation in the worst-case setting},
%    {arXiv:1909.02290v1}.
%  %{2019}.
 
%  \bibitem{Kam_Vol19} L. K\"ammerer and T.  Volkmer, 
% {\it Approximation of multivariate periodic functions based on sampling along multiple rank-1 lattices},
%  {J. Approx. Theory},
%   {246}, {1--27},
%   {2019},
% https://doi.org/10.1016/j.jat.2019.05.001.

% \bibitem{KieWo60}
% J.~Kiefer and J.~Wolfowitz.
% \newblock {\it The equivalence of two extremum problems}, 
% \newblock Canadian J. Math. 12, 363--366, 1960.


% \bibitem{Kosov_Temlyakov23} 
% E.\,D. Kosov and V.\,N. Temlyakov,
% {\it Sampling discretization of the uniform norm and applications},
% arXiv:2306.14207, 2023.

\bibitem{KNU24}
 D. Krieg, E. Novak, M. Ullrich, 
\textit{On the power of adaption and randomization}, 
arXiv:2406.07108. 

% \bibitem{KPUU23}
% D.\,Krieg, K.\,Pozharska, M.\,Ullrich, T.\,Ullrich,
% {\it Sampling recovery in $L_2$ and other norms},
% arXiv:2305.07539, 2023.


% \bibitem{KSUW} D. Krieg, P. Siedlecki, M. Ullrich, and H. Wo\'zniakowski,
% {\it Exponential tractability of $L_2$-approximation with function values},
% Adv. Comput. Math. 49, 18, 2023. %https://doi.org/10.1007/s10444-023-10021-7

% \bibitem{KSmani} D. Krieg, M. Sonnleitner,
% {\it Function recovery on manifolds using scattered data}, arXiv:2109.04106, 2021.


% \bibitem{KU1} D. Krieg and M. Ullrich,
% {\it Function values are enough for $L_2$-approximation}, 
% Found. Comp. Math., 21(4), 1141--1151, 2021.

% \bibitem{KU2} D.\,Krieg and M.\,Ullrich,
% {\it Function values are enough for $L_2$-approximation: Part II}, 
% J. Complexity 66, 101569, 2021. % https://doi.org/10.1016/j.jco.2021.101569.

% \bibitem{KV} D.\,Krieg and J.\,Vyb\'iral, {\it New lower bounds for the integration of periodic functions}, arXiv:2302.02639, 2023.

% \bibitem{Kunsch2018}
%    R.\,Kunsch,
%  {\it Breaking the curse for uniform approximation in {H}ilbert spaces via {M}onte {C}arlo methods},
%   {J. Complexity},
%   {48},  {15--35},
%   {2018},
% https://doi.org/10.1016/j.jco.2018.04.002.


% \bibitem{KSU} T. Kühn, W. Sickel, and T. Ullrich, {\it Approximation of mixed order Sobolev functions on the $d$-torus -- asymptotics, preasymptotics and $d$-dependence}.
% Constructive Approximation, 42, 353--398, 2015.


%\bibitem{KWW08}
%F.\,Kuo, G.\,W.\,Wasilkowski, and H.\,Wo\'zniakowski,
%{\it Multivariate {$L_\infty$} approximation in the worst case setting over reproducing kernel {H}ilbert %spaces},
%Journal of Approximation Theory, 152(2), 135--160, 2008,
%https://doi.org/10.1016/j.jat.2007.11.006.


%\bibitem{KWW}
%F.~Kuo, G.~W.~Wasilkowski, and H.~Wo\'zniakowski,
%{\it On the power of standard information for multivariate approximation
%  in the worst case setting},
%Journal of Approximation Theory, 158, 97--125, 2009.

%\bibitem{LS} Y. T. Lee and H. Sun, {\it Constructing linear-sized spectral sparsification in almost-linear time}, SIAM Journal on Computing, 47(6), 2315-2336, 2018.

\bibitem{LE11} J. Lang and D.E. Edmunds. \textit{Eigenvalues, embeddings and generalised trigonometric functions}, volume
2016 of Lecture Notes in Mathematics, Springer, Heidelberg, 2011.

% \bibitem{LS} 
% J. Leveslay and X. Sun, {\it Approximation in rough native space by shifts of smooth kernels on spheres}, J. Approx. Theory
% 133, 269--283, 2005.

% \bibitem{LT} I. Limonova and V. Temlyakov, {\it On sampling discretization in $L_2$}, J. Math. Anal. Appl. 515(2), 126457, 2022. 


%\bibitem{LW} W. Lu and H. Wang, {\it On the power of standard information for tractability for $L_2$-approximation in the randomized setting}, Contemporary Mathematics, 3(1), 1--29, 2022.

%\bibitem{MSS} A. Marcus, D. Spielman and N. Srivastava, {\it Interlacing families II: Mixed characteristic polynomials and the Kadison-Singer problem}, Annals of Mathemathics, 327-350, 2015, http://dx.doi.org/10.4007/annals.2015.182.1.8.

% \bibitem{Maiorov1976}
% V.\,E. Maiorov,  {\it The best approximation of classes $W^r_1(I^s)$  in the space $L_\infty(I^s)$}, 
% Mat. Zametki 19(5), 699--706, 1976.

\bibitem{Mathe90}
P. Math\'e, \textit{s-Numbers in Information-Based Complexity}, J. Complexity 6, 41--66, 1990.

%\bibitem{MV92} R. Meise and D. Vogt, {\it Einf\"uhrung in die Funktionalanalysis},
%\newblock Braunschweig; Wiesbaden: Vieweg, 1992 (Vieweg-Studium; 62: Aufbaukurs Mathematik).


%\bibitem{MP} S. Mendelson and A. Pajor, {\it On singular values of matrices with independent rows}, Bernoulli, 12(5), 761-773, 2006.

%\bibitem{Mit} B.\,S.\,Mityagin, {\it Approximation of functions in $L^p$ and $C$ on the torus}.
%Mathematical Notes, 58, 397--414, 1962.

\bibitem{MH63}
B. S. Mityagin, G. M. Henkin, 
\textit{Inequalities between n-diameters}, in Proc. of the Seminar on Functional Analysis 7,
Voronezh, 97--103, 1963.

% \bibitem{NSU} N. Nagel, M. Sch\"afer, and T. Ullrich,
% {\it A new upper bound for sampling numbers}, Found. Comp. Math. 22(2), 445-468, 2022.

% \bibitem{Nevai} P. Nevai. 
% \emph{G\'eza Freud, orthogonal polynomials and Christoffel functions. A case study. } J.
% Approx. Theory, 48(1):3--167, 1986.

%\bibitem{NOU2011} S. Nitzan, A. Olevskii and A. Ulanovskii, {\it A few remarks on sampling of signals with small spectrum}, Proceedings of the Steklov Institute of Mathematics, 280(1), 240-247, 2013.

%\bibitem{NOU2016} S. Nitzan, A. Olevskii and A. Ulanovskii, {\it Exponential frames on unbounded sets}, Proceedings of the American Mathematical Society, vol. 144, no 1, 109-118, 2016, http://dx.doi.org/10.1090/proc/12868.

% \bibitem{NNS} V.\,D.\,Nguyen, V.\,K.\,Nguyen, and W.\,Sickel, {\it s-numbers of embeddings of weighted
% Wiener algebras}. J. Approx. Theory, 279, Paper No. 105745, 2022.


\bibitem{Novak} E. Novak, {\it Deterministic and stochastic error bounds in numerical analysis},
Lecture Notes in Mathematics 1349, Springer-Verlag, 1988.

%\bibitem{NoTr06}
%E.~Novak and H.~Triebel, 
%\emph{Function Spaces in Lipschitz Domains and
%Optimal Rates of Convergence for Sampling}, 
%Constr. Approx. (2006) 23: 325–350.



% \bibitem{NW1}
% E.~Novak and H.~Wo\'zniakowski,
% \newblock {\em Tractability of multivariate problems. {V}olume~{I}: {L}inear
%   information}, volume~6 of {\em EMS Tracts in Mathematics}.
% \newblock European Mathematical Society (EMS), Z\"urich, 2008.

%\bibitem{NW2}
%E.~Novak and H.~Wo\'zniakowski.
%\newblock {\em Tractability of multivariate problems. {V}olume {II}: {S}tandard
%  information for functionals}, volume~12 of {\em EMS Tracts in Mathematics}.
%\newblock European Mathematical Society (EMS), Z\"urich, 2010.

% \bibitem{NW3} E.\,Novak and H.\,Wo\'zniakowski, {\it Tractability of multivariate problems. Volume~III: Standard information for operators},
% volume~18 of {\em EMS Tracts in Mathematics}.
% \newblock European Mathematical Society (EMS), Z\"urich, 2012.

%\bibitem{NW16}
%E.~Novak and H.~Wo\'zniakowski.
%\newblock {\it Tractability of multivariate problems for standard and linear
%  information in the worst case setting: part I},
%\newblock {Journal of Approximation Theory}, 207, 177--192, 2016.


% \bibitem{Osipenko_Parfenov1995}
% K.\,Y. Osipenko and  O.\,G.\,Parfenov, 
% {\it Ismagilov type theorems for linear, {G}el'fand and {B}ernstein
%           {$n$}-widths},
%  {J. Complexity},
%  {11(4)}, {474--492},
%  {1995},
% http://dx.doi.org/10.1006/jcom.1995.1025.

% \bibitem{PePe}
% K. B. Petersen and M. S. Pedersen,
% {\it The Matrix Cookbook},
% Technical University of Denmark, Version 20121115, 2012, 
% http://www2.compute.dtu.dk/pubdb/pubs/3274-full.html

\bibitem{Pietsch-s}
A. Pietsch, {\it s-Numbers of operators in Banach spaces}, 
Studia Math. 51, 201--223, 1974.

% \bibitem{Pietsch-ideals} A. Pietsch. {\it Operator ideals}, 
% % Deutscher Verlag, Wiss., Berlin, 1978.
% North-Holland Mathematical Library 20, Elsevier, 1980.

\bibitem{Pietsch-Weyl}
A. Pietsch, 
{\it Weyl  Numbers  and  Eigenvalues  of  Operators  in  Banach Spaces}, 
Math. Ann. 247,  I49--168, 1980. 

\bibitem{Pie87}
A. Pietsch. 
{\it Eigenvalues and s-numbers}, 
Cambridge studies in advanced mathematics 13, Cambridge University Press, 1987.

\bibitem{Pie07} A. Pietsch, {\it History of Banach spaces and linear operators}, Birkhäuser Boston, MA, 2007.

\bibitem{Pie08}
A. Pietsch, 
{\it Bad properties of the Bernstein numbers}, 
Studia Math. 184 (3), 263--269, 2008.

\bibitem{Pie09}
A. Pietsch, \textit{Long-standing open problems of Banach space
theory: My personal top ten,} Quaestiones Mathematicae 32:3, 321--337, 2009. 
% DOI: 10.2989/QM.2009.32.3.4.905

\bibitem{Pinkus85}
A. Pinkus, {\it n-width in approximation theory}, 
Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge / A Series of Modern Surveys in Mathematics, %(MATHE3, volume 7), 
Springer Berlin, Heidelberg,~1985. 
%https://doi.org/10.1007/978-3-642-69894-1.


% \bibitem{PU}
% K. Pozharska and T. Ullrich.
% {\it A note on sampling recovery of multivariate functions in the uniform norm}, 
% SIAM J. Numer. Anal., 60(3),  1363--1384, 2022. 
% %https://doi.org/10.1137/21M1410580. (arXiv:2103.11124v2)

% \bibitem{Rud99}
% M. Rudelson, 
% {\it Almost orthogonal submatrices of an orthogonal matrix}, Isr. J. Math. 111, 143--155, 1999.

%\bibitem{S41}
%R. Salem. 
%{\it A new proof of a theorem of Menchoff},
%Duke Mathematical Journal, 8(2), 269--272, 1941.

\bibitem{Schmidt1907}
E. Schmidt,  
\textit{Entwicklung willk\"urlicher Funktionen nach Systemen vorgeschriebener},  Math.  Ann.  63, 
433-476, 1907.

%\bibitem{SU}       
%W. Sickel and T. Ullrich,  
%{\it Smolyak's algorithm, sampling on sparse grids and function spaces        
%of dominating mixed smoothness},        
%East Journal on Approximations, 13(4), 387--425, 2007.  


% \bibitem{Schmidt1944} 
%  E. Schmidt, 
% {\it \"Uber die nebst ihren Ableitungen orthogonalen Polynomensysteme und das zugeh\"orige Extremum},
% Math. Ann., 119, 165--204, 1944, https://doi.org/10.1007/BF01563739.

% \bibitem{Schn14} R. Schneider, {\it Convex bodies: the Brunn-Minkowski theory}, 2nd expanded ed. (English),
% Encyclopedia of Mathematics and its Applications 151. Cambridge University Press, 2014. 
% % (ISBN 978-1-107-60101-7/hbk; 978-1-139-00385-8/ebook). xxii, 736 p. 

% \bibitem{Sh95}
% B.~Shekhtman, 
% {\it Interpolation of individual functions}, 
% Comput. Math. Appl. 30, 191--196, 1995.

\bibitem{SX24}
J.W. Siegel and J. Xu. \textit{Sharp bounds on the approximation rates, metric entropy, and n-widths of shallow
neural networks}, Found. Comput. Math. 24(2):481–537, 2024.

% \bibitem{Sm}        
% S. A. Smolyak,       
% {\it Quadrature and interpolation formulas for tensor products       
% of certain classes of functions},     
% Soviet Mathematics Doklady, 4, 240--243, 1963. 

% \bibitem{Steinwart_Scovel2012}
% I. Steinwart and C. Scovel,
% {\it Mercer’s theorem on general domains:
% on the interaction between measures, kernels,
% and {RKHS}s},
% Constr. Approx., 35, 363--417, 2012, 
% https://doi.org/10.1007/s00365-012-9153-3.

%\bibitem{Szego1975} 
% G. Szeg\"{o},
%{\it Orthogonal Polynomials},
%4th edition Amer. Math. Soc., Providence, 1975.


%\bibitem{Tem93}
%V.\,N.\,Temlyakov, 
%{\it On approximate recovery of functions with bounded mixed derivative}, J.
% Complexity, 9,  41--59, 1993, https://doi.org/10.1006/jcom.1993.1004. 

%V.~N. Temlyakov.
%\newblock {\em Approximation of periodic functions}, 
%	Computational Mathematics and Analysis Series, 
%	Nova Science Publishers, Inc., Commack, NY, 1993.

%\bibitem{Tem18}
%V.~N. Temlyakov.
%\newblock {\em Multivariate Approximation}, volume~32 of {\em Cambridge
%  Monographs on Applied and Computational Mathematics}.
%\newblock Cambridge University Press, 2018.

% \bibitem{T20} V.\,N.\,Temlyakov, {\it On optimal recovery in $L_2$},  
% J. Complexity, 65, 101545, 2021. % https://doi.org/10.1016/j.jco.2020.101545.

% \bibitem{T23} V.\,N.\,Temlyakov, {\it 
% On universal sampling recovery in the uniform norm}, arXiv:2307.04017.


%\bibitem{T18} V. N. Temlyakov, 
%{\it The Marcinkiewicz-type discretization theorems}, 
%Constructive Approximation, 48, 337-369, 2018.

% \bibitem{TU1}
% V.~N.~Temlyakov and T.~Ullrich. 
% {\it Approximation of functions with small mixed smoothness in the uniform norm},
% Journal of Approximation Theory, 277, 105718, 2022.

%\bibitem{TU2} V. N. Temlyakov and T. Ullrich, 
%{\it Bounds on Kolmogorov widths and sampling recovery for
%classes with small mixed smoothness}, Journal of Complexity, 67, 101575, 2021.

\bibitem{Tikh60} 
V. M. Tikhomirov, \textit{Diameters of sets in
function spaces and the theory of best approximations}, 
Russ. Math. Survey 15(3), 75--111, 1960. 

%\bibitem{TWW88}
%J.~Traub, G.~Wasilkowski, and H.~Wo\'zniakowski, \emph{Information-Based Complexity}, 
%Acad.~Press, 1988.

%\bibitem{Tr11} J. Tropp, {\it User-friendly tail bounds for sums of random matrices}, Found. Comp. Math., %12(4):389–434, 2011.

%\bibitem{U2020} M. Ullrich, {\it On the worst-case error of least squares algorithms for $L_2$-approximation %with high probability}, J.~Complexity, 60, 2020.


%\bibitem{V} J. Vyb\'\i ral, 
%{\it A variant of Schur's product theorem and its applications},
%Advances in Mathematics, 368, 107140, 2020.

% \bibitem{Vyb-widths}
% J. Vyb\'iral, \textit{Widths of embeddings in function spaces}, J. Complexity, 24(4), 545--570, 2008.

%\bibitem{WW95}         
%G. W. Wasilkowski and H. Wo\'zniakowski,       
%{\it Explicit cost bounds of algorithms for multivariate  tensor product problems},
%Journal of Complexity, 11, 1--56, 1995.

%\bibitem{WW01}
%G.\,W.~Wasilkowski and H.~Wo\'zniakowski. 
%{\it On the power of standard information for weighted approximation},
%Foundations of Computational Mathematics, 1(4), 417-434, 2001.

%\bibitem{WW07} G.W. Wasilkowski and H. Wo\'zniakowski, 
%{\it The power of standard information for multivariate approximation 
%in the randomized setting}, Mathematics of Computation, 76, 965-988, 2007.

\bibitem{Weyl1949}
H. Weyl, 
\textit{Inequalities between the two kinds of eigenvalues of a linear transformation}, Proc.
Nat. Acad. Sci. USA 35, 408--411, 1949.


% \bibitem{XN23}
% Y.~Xu, A.~Narayan, 
% \emph{Randomized weakly admissible meshes}, 
% J.~Approx. Theory 285, 105835, 2023.


\end{thebibliography}

\documentclass{birkjour}

\usepackage[utf8]{inputenc}
\usepackage{amsmath, amsthm, amsfonts, amssymb}
\usepackage{hyperref} 
\usepackage{enumitem}

\newtheorem{theorem}{Theorem}

\newtheorem{coro}[theorem]{Corollary}
\newtheorem{prop}[theorem]{Proposition}
\theoremstyle{definition}
\newtheorem{remark}[theorem]{Remark}

\newcommand{\eps}{\varepsilon}
\renewcommand{\phi}{\varphi}

\newcommand{\N}{\mathbb{N}}

\newcommand{\ipr}[1]{\left\langle#1\right\rangle}
\newcommand{\set}[1]{\Bigl\{#1\Bigr\}}
\newcommand{\bset}[1]{\Biggl\{#1\Biggr\}}
\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\co}{\colon}

\newcommand{\calL}{\mathcal{L}}

\DeclareMathOperator{\Span}{span}
\DeclareMathOperator{\rank}{rank}
\DeclareMathOperator{\codim}{codim}

\begin{document}

%
%
%
%
%
%
%
%
%

\title[Inequalities between s-numbers]{Inequalities between s-numbers}

\author{
Mario Ullrich}
\address{Institut f\"ur Analysis, 
Johannes Kepler Universit\"at Linz, Austria
} 
\email{mario.ullrich@jku.at}

\subjclass{
Primary 47B06;      
Secondary 
46B50,  
47B01
}

\keywords{s-numbers, Hilbert numbers, Pietsch}

\date{September 16, 2014}


\dedicatory{In memory of Albrecht Pietsch}


\begin{abstract} 
Singular numbers of linear operators between Hilbert spaces were generalized to Banach spaces by s-numbers (in the sense of Pietsch). 
This allows for different choices, including 
approximation, Gelfand, Kolmogorov and Bernstein numbers. 
Here, we present an elementary proof of a bound between the smallest and the largest s-number. 
\end{abstract}

\maketitle

\medskip

We start with introducing the terminology and a presentation of the results. 
In Section~\ref{sec:hist} we will discuss them and some
history.  
Proofs are given in Section~\ref{sec:proof}.

\section{s-numbers}

In what follows, let $X$, $Y$, $Z$ and $W$ be real or complex Banach spaces.  
The (closed) unit ball of $X$ is denoted by $B_X$, 
the dual space of $X$ by $X'$, and the identity map on $X$ is denoted by $I_X$. 
For a closed subspace $M\subset X$, we write $J_M^X$ for the embedding 
$J_M^X\co M\to X$ with $J_M^X(x)=x$, 
and $Q_M^X$ for the canonical map $Q_M^X\co X\to X/M$ with $Q_M^X(x)=x+M$
onto the quotient space 
$X/M:=\{x+M\co x\in X\}$ with norm $\|x+M\|_{X/M}:=\inf_{m\in M}\|x+m\|_X$.
The dimension of a subspace $M\subset X$ is denoted by $\dim(M)$, and by $\codim(M):=\dim(X/M)$ we denote its codimension.

The class of all bounded linear operators between Banach spaces 
is denoted by $\calL$, and by $\calL(X,Y)$ we denote those operators from $X$ to $Y$, equipped with the operator norm. 
The rank of an operator $S\in\calL(X,Y)$ is defined by $\rank(S):=\dim(S(X))$.

\medskip
\goodbreak

A map $S\to (s_n(S))_{n\in\N}$ assigning to every operator $S\in\calL$ a non-negative scalar sequence $(s_n(S))_{n\in\N}$ is called an 
\textit{s-number sequence} if, for all $n\in\N$, the following conditions are satisﬁed

\begin{enumerate}[label=(S\arabic*)]
\itemsep=2mm
\item $\|S\|=s_1(S)\ge s_2(S) \ge \dots \ge 0$ \qquad for all \; $S\in\calL$,
\item $s_n(S+T) \,\le\, s_n(S)+\|T\|$ \qquad for all \; $S,T\in\calL(X,Y)$,
\item $s_n(BSA) \,\le\, \|B\| \, s_n(S) \, \|A\|$ \qquad where \;$W \stackrel{A}{\longrightarrow} X\stackrel{S}{\longrightarrow} Y\stackrel{B}{\longrightarrow} Z$, 
\item $s_n(I_{\ell_2^n})=1$, 
\item $s_n(S)=0$ \quad whenever $\rank(S)<n$.
\end{enumerate}
\medskip
We call $s_n(S)$ the \textit{$n^{\rm th}$ s-number} of the operator $S$. To indicate the underlying Banach spaces, we sometimes write $s_n(S\co X\to Y)$.

\goodbreak

There are some especially important examples of s-numbers: 
\begin{itemize}
    \item approximation numbers: 
    \[
    a_n(S) \,:=\, \inf\set{\|S-L\|\co L\in\calL(X,Y), \; \rank(L)<n}
    \]
    \item Bernstein numbers: 
    \[
    b_n(S) \,:=\, \sup\set{\inf_{x\in M\setminus\{0\}}\frac{\|Sx\|}{\|x\|}\co M\subset X, \; \dim(M)=n}
    \]
    \item Gelfand numbers: 
    \[
    c_n(S) \,:=\, \inf\set{\|S J_M^X\|\co M\subset X, \; \codim(M)<n} 
    \] 
    \item Kolmogorov numbers: 
    \[
    d_n(S) \,:=\, \inf\set{\|Q_N^Y S\|\co N\subset Y, \; \dim(N)<n}
    \]
    \item Weyl numbers: \\[-4mm]
    \[
    x_n(S) \,:=\, \sup\bset{\frac{a_n(SA)}{\|A\|}\co A\in\calL(\ell_2,X),\; A\neq 0}
    \]
    \item Hilbert numbers: 
    \[
    \begin{split}
    h_n(S) \,:&=\, \sup\bset{\frac{a_n(BSA)}{\|B\|\|A\|}\co A\in\calL(\ell_2,X), \; B\in\calL(Y,\ell_2),\; A,B\neq0}. \\
    \end{split}
    \]
\end{itemize}

\bigskip 

We refer to~\cite{Pie87,Pie07} for a detailed treatment of the above, and a few other, s-numbers and their specific properties.  

\begin{remark} \label{rem:def}
The original definition of s-numbers in~\cite{Pietsch-s} 
used the stronger \textit{norming property} 
$(S4')\co s_n(I_X)=1$ for all $X$ with $\dim(X)\ge n$. 
This did not allow for $x_n$ and $h_n$, 
and has been weakened in~\cite{Bauhardt,Pietsch-Weyl} for defining 
them, leading to the least restrictive axioms that still imply uniqueness for Hilbert spaces, see Proposition~\ref{prop1}.  
It is sometimes assumed that the s-number sequence is \textit{additive}, i.e., 
$(S2')\co s_{m+n-1}(S+T)\le s_m(S)+s_n(T)$, 
or \textit{multiplicative}, i.e., 
$(S3')\co s_{m+n-1}(ST)\le s_m(S) s_n(T)$. 
Both properties hold for $a_n,c_n,d_n,x_n$, 
while $h_n$ is only additive, 
and $b_n$ is neither of the two, 
see~\cite{Pie07,Pie08}. 
\end{remark}

\goodbreak 

The following proposition is well-known, see~\cite[2.3.4 \& 2.6.3 \& 2.11.9]{Pie87}. 

\begin{prop} \label{prop1}
For every s-number sequence $(s_n)$, $S\in\calL$ and $n\in\N$, we have 
\[
h_n(S) \,\le\, s_n(S) \,\le\, a_n(S). 
\]
Equalities hold if $S\in\calL(H,K)$ for Hilbert spaces $H$ and $K$.
\end{prop}


For convenience, 
we present a sketch of the proof of the inequalities in Section~\ref{sec:proof}. 
Using only elementary arguments, 
we prove the following reverse inequality,  
which is known in a 
more involved form 
based on an intermediate comparison with $x_n(S)$, see~\cite[2.10.7]{Pie87} or~\cite[6.2.3.14]{Pie07}, 
or Remark~6 in~\cite{KNU24}.


\begin{theorem}\label{thm:c-h}
For all $S\in\calL$ and $n\in\N$, we have 
\[
\max\set{c_n(S),\, d_n(S)} 
\;\le\; n\,\left(\prod_{k=1}^n h_k(S) \right)^{1/n}.
\]    
\end{theorem}

\medskip

Since $c_n(I\co \ell_1\to\ell_\infty)\asymp 1$ 
and $h_n(I\co \ell_1\to\ell_\infty)\asymp n^{-1}$, 
see~\cite[6.2.3.14]{Pie07} and~\cite{HL84}, 
this result is best possible up to constants.

\bigskip 
\goodbreak

We cannot obtain bounds for individual $n$ from Theorem~\ref{thm:c-h}, 
see also Remark~\ref{rem:eig}, 
but combining it with the inequality 
$n^\alpha\le e^\alpha\, (n!)^{\alpha/n}$ 
for $\alpha\ge0$, 
we obtain a 
more handy form. 
Moreover, the known fact that 
$a_n(S)\le (1+\sqrt{n})\, c_n(S)$, 
see~
\cite[2.10.2]{Pie87}, leads to a bound between $a_n$ and $h_n$, and hence between all s-numbers.  

\medskip

\begin{coro}
For all $S\in\calL$, $\alpha>0$ and $n\in\N$, we have 
\[
c_n(S) \,\le\, e^\alpha\, n^{-\alpha+1}\cdot 
\sup_{k\le n}\, k^\alpha\, h_k(S),
\]
and
\[
a_n(S) \,\le\, 2\,e^\alpha\, n^{-\alpha+3/2}\cdot 
\sup_{k\le n}\, k^\alpha\, h_k(S).
\]
\end{coro}





\section{A bit of history} \label{sec:hist}

We provide a brief description 
of the relevant facts from 
(the highly recommended) 
``History of Banach Spaces
and Linear Operators'' of Pietsch~\cite{Pie07}, 
and give some further references.

\textbf{Singular numbers} 
of operators 
on Hilbert spaces 
have become 
fundamental 
tools in (applied) mathematics 
since their introduction in 1907 by Schmidt~\cite{Schmidt1907}. 
For compact $S\in\calL(H,K)$ between complex Hilbert spaces~$H,K$, 
the \textit{singular numbers} are 
defined by 
$s_k(S):=\sqrt{\lambda_k(SS^*)}$, 
where the \textit{eigenvalues} $\lambda_k(T)$ of $T\in\calL(X,X)$ are characterized by $Te_k=\lambda_k(T)\cdot e_k$ for some $e_k\in X\setminus\{0\}$, and ordered decreasingly. 
%
Applications range from 
the study of \textit{eigenvalue distributions} of operators,  
see~\cite{Koe86} or~\cite[6.4]{Pie07}, 
to the classification of \textit{operator ideals}~\cite[6.3]{Pie07}, 
to the singular value decomposition, aka Schmidt representation, 
with its many applications.

\textbf{\!\!\!s-numbers} are a generalization to 
linear operators between Banach spaces. 
However, 
there is no unique substitute for singular numbers but, 
depending on the context, different 
s-numbers may be used to 
\textit{quantify compactness}, 
while 
others may be easier to compute. 
As Pietsch wrote in~\cite[6.2.2.1]{Pie07}, 
``we have a large variety of s-numbers that make our life more interesting.''

Most notably, 
$a_n$, $b_n$, $c_n$ and $d_n$ were already known in the 1960s, 
sometimes in a related form as ``width'' of a set, 
see~\cite{Ism74,Tikh60} or~\cite[6.2.6]{Pie07}, 
and are 
by now part of the foundation of approximation theory~\cite{Pinkus85} 
and information-based complexity~\cite{Mathe90,Novak}. 
More recent treatments of the subject 
and extensions 
can be found, e.g., in~\cite{FMS21,LE11,SX24}. 
Let us also highlight~\cite{KNU24}, where 
we discuss the relation of s-numbers to \textit{minimal approximation errors} in detail, 
and use variants of Theorem~\ref{thm:c-h} 
to bound the \emph{maximal gain} of randomized/adaptive algorithms over 
deterministic/non-adaptive ones. 

An axiomatic theory of s-numbers has been developed by Pietsch~\cite{Pietsch-s,Pie87,Pie07} in the 1970s. 
This, in particular, allowed for a characterization of the smallest/largest s-number (with certain properties), 
see also Remark~\ref{rem:def}. 

\textbf{Inequalities} 
and several relations between s-numbers have already been collected 
in~\cite{Pietsch-s}, see also~\cite[2.10]{Pie87} and~\cite[6.2.3.14]{Pie07}. 
A particularly interesting bound is 
$d_n(S)\le n^2 b_n(S)$, which was proven 
by Mityagin and Henkin~\cite{MH63} in 1963. 
They also conjectured that $n^2$ can be replaced by $n$, see also~\cite[p.~24]{Pinkus85}. 
This bound with $b_n$ replaced by $h_n$, 
and the corresponding conjecture, have been given in~\cite{Bauhardt}, 
and apparently, the bound has not been improved since then. 
However, in the weaker form as in Theorem~\ref{thm:c-h}, this problem has been solved by Pietsch~\cite{Pietsch-Weyl} in 1980, see the final remarks there. In particular, it is shown that 
there is some $C_\alpha>0$ such that 
$\sup_{k\ge1} k^\alpha d_k(S) \le C_\alpha\, \sup_{k\ge1} k^{\alpha+1} h_k(S)$, see also~\cite[2.10.7]{Pie87} or~\cite[6.2.3.14]{Pie07}. 

\goodbreak

\textbf{The proof} 
of this result is the blueprint for the proof of 
Theorem~\ref{thm:c-h}. 
However, those proofs employ an intermediate comparison with the 
Weyl numbers $x_n$, which lie somehow 
between 
$h_n$ and $c_n$. 
Despite interesting consequences, 
this approach requires the multiplicativity of $x_n$, the notion of 2-summing norm, and some technical difficulties. 
The proof presented here is elementary: It 
only uses 
definitions and 
known properties of the determinant.

\textbf{Bounds for individual $n$} 
cannot be deduced from this approach, see also Remark~\ref{rem:eig}, 
and it remains a long-standing open problem if 
$d_{bn}(S)\le c\, n\, b_n(S)$, 
or even 
$a_{bn}(S)\le c\, n\, h_n(S)$, for some $b,c\ge1$, see~\cite[Prob.~5]{Pie09} or~\cite[2.10.7]{Pie87}.

\goodbreak 


\section{The proofs} \label{sec:proof}

\begin{proof}[Proof of Proposition~\ref{prop1}]
We refer to~\cite[2.11.9]{Pie87} for the proof that 
$s_n(S)=a_n(S)$ for any s-number sequence $(s_n)$, and any $S\in\calL(H,K)$ for Hilbert spaces~$H$ and~$K$. 
Just note that, for compact $S$, this follows quite directly from the singular value decomposition.
From this and (S3), we obtain 
\[
h_n(S) 
\,=\, \sup\bset{\frac{s_n(BSA)}{\|B\|\|A\|}\co A\in\calL(\ell_2,X), \; B\in\calL(Y,\ell_2), \; A,B\neq 0} 
\,\le\, s_n(S)
\] 
for any $(s_n)$. 
In addition, 
by (S2) and (S5), 
we obtain for any $L$ with $\rank(L)<n$ that
$
s_n(S)
\,\le\, s_n(L) + \|S-L\| 
\,=\, \|S-L\| 
$.
By taking the infimum over all such $L$, we see that 
$s_n(S)\le a_n(S)$. 
\end{proof}

\medskip

\begin{proof}[Proof of Theorem~\ref{thm:c-h}]
We first present the proof from~\cite[2.10.3]{Pie87} of the following statement: 
For fixed $\eps>0$, we can find 
$x_1,\dots,x_n\in B_X$ and $b_1,\dots,b_n\in B_{Y'}$ 
such that $\ipr{Sx_k,b_j}=0$ for $j<k$ and 
$(1+\eps)|\ipr{Sx_k,b_k}|>c_k(S)$ for $k=1,\ldots,n$. 

For this, we inductively assume that $x_k,b_k$ for $k<n$ are already found, and define 
\[
M_n \,:=\, \set{x\in X\co \ipr{Sx,b_k}=0 \text{ for } k<n}. 
\] 
Since $\codim{M_n}<n$, we can choose $x_n\in M_n \cap B_X$ 
with 
\[
(1+\eps) \|Sx_n\| \,\ge\, \|S J_{M_n}^X\| \,\ge\, c_n(S).  
\]
By the Hahn-Banach theorem, we choose $b_n\in B_{Y'}$ with $\ipr{Sx_n,b_n}=\|Sx_n\|\ge\frac{c_k(S)}{1+\eps}$. 

We now define the operators 
\[
A(\xi) \,:=\, \sum_{i=1}^n \xi_i x_i\in X, \quad \xi=(\xi_i)\in\ell_2^n,
\]
and
\[
B(y) \,:=\, \bigl(\ipr{y,b_i}\bigr)_{i=1}^n \in\ell_2^n, \qquad y\in Y, 
\medskip
\]
which satisfy $\|A\|,\|B\|\le \sqrt{n}$, 
and observe that $S_n:=BSA\co\ell_2^n\to\ell_2^n$ is generated by the triangular matrix $(\ipr{Sx_i,b_j})_{i,j=1}^n$ 
with determinant $\det(S_n)\ge 
\prod_{k=1}^n \frac{c_k(S)}{1+\eps}$. 

To obtain a bound with s-numbers, note that 
they all coincide for~$S_n$, esp.~with $a_k(S_n)$, and are equal to the 
singular numbers $s_k(S_n)$, i.e., the roots of the eigenvalues of $S_n S_n^*$, 
see~\cite[6.2.1.2]{Pie07}. 
As the determinant is multiplicative and equals the product of the eigenvalues, we see that 
$\det(S_n)=\sqrt{\det(S_nS_n^*)}=\prod_{k=1}^n a_k(S_n)$. 
%
From the definition of $h_n$, we obtain 
$a_k(S_n)\le \|A\|\|B\|\,h_k(S)\le n\cdot h_k(S)$, 
and hence 
\[\begin{split}
(1+\eps)^{-n}\prod_{k=1}^n c_k(S)  
\;&\le\;\det(S_n) 
\;=\;  \prod_{k=1}^n a_k(S_n) 
\;\le\;  n^n \prod_{k=1}^n h_k(S). 
\end{split}\]
With $\eps\to0$ and $c_n(S)\le \bigl(\prod_{k=1}^n c_k(S)\bigr)^{1/n}$ we obtain the result for $c_n(S)$. 

\goodbreak

The proof for $d_n(S)$ could be done via duality, 
at least for compact $S$, 
see e.g.~\cite[6.2.3.9 \& 6.2.3.12]{Pie07}. 
However, one can also prove it directly 
by inductively choosing $M_n:=\Span\{Sx_k\co k<n\}$, 
$x_n\in B_X$ with 
$(1+\eps) \|Q^Y_{M_n}Sx_n\| \,\ge\, \|Q^Y_{M_n}S\| \,\ge\, d_n(S)$, 
and $b_n\in B_{Y'}$ with 
$\ipr{Sx_n,b_n}=\|Q^Y_{M_n}Sx_n\|$ 
and  
$\ipr{Sx_k,b_n}=0$ for $k<n$. 
The remaining proof is as above. 
\end{proof}

\goodbreak

\begin{remark} \label{rem:eig}
The proof of Theorem~\ref{thm:c-h} uses the determinant 
to relate the eigenvalues $\lambda_k(S_n)$ of $S_n$ (which are the diagonal entries) with its singular numbers. 
Sometimes, the more general \textit{Weyl's inequality}~\cite{Weyl1949} from 1949 is used, which states that 
$
\prod_{k=1}^n |\lambda_k(S)| 
\;\le\; \prod_{k=1}^n a_k(S)
$
for any 
compact $S\in\calL(H,H)$, 
see also~\cite[3.5.1]{Pie07}.
This crucial step appears in all the proofs I am aware of 
that lead to the optimal factor~$n$ in the comparisons. 
Unfortunately, all these approaches use a whole collection of s-numbers, which does not allow for bounds for individual $n$. 

Let us present an example from~\cite[2.d.5]{Koe86} that shows that 
such product bounds 
between eigenvalues and s-numbers 
are to some extent best possible:\\
For $0<\sigma<1$, consider the matrix 
$T_n=(\delta_{j,i+1}+\sigma\cdot \delta_{i,n}\delta_{j,1})_{i,j=1}^n$ 
which represents a mapping on $\ell_2^n$. 
It is easy to verify that $a_k(T_n)=1$ for $k<n$ and $a_n(T_n)=\sigma$. 
(Recall that $a_k$ are the singular numbers in this case.)
Moreover, since $T_n^n=\sigma\cdot I_{\ell_2^n}$, 
we see that $|\lambda_k(T_n)|=\sigma^{1/n}$ for $k=1,\dots,n$. 
This shows that Weyl's inequality, 
as well as the easy corollary 
$|\lambda_n(S)|\le \|S\|^{1-\frac1n} a_n(S)^{1/n}$, 
are in general best possible. 
\end{remark}

\smallskip

\noindent 
{\bf Acknowledgement: } 
I thank Albrecht Pietsch for comments on an earlier version, and for encouraging me to make this note 
as self-contained as it is now. 
The last comments I received from him were on March 8, 2024, where he added that he was busy with his own article. 
My reply, however, was answered by his granddaughter, who informed me that Professor Pietsch had passed away on March 10. \\
I also thank S. Heinrich, A.~Hinrichs, D.~Krieg, T. K\"uhn, E.~Novak and the anonymous referees for helpful comments. 



\end{document}